# Gubkin str. 8, 119991 Moscow, Russia

### v1

@@ 2 Setup: theory and its decomposition
The resolution for $(\omega^{2}\mathbf{I}+|\mathbf{p}|^{2}\mathbf{S}(\hat{\mathbf{n}}))^{-1}$ then reads
$(\omega^{2}\mathbf{I}+|\mathbf{p}|^{2}\mathbf{S}(\hat{\mathbf{n}}))^{-1}=\sum_{a=1}^{2}\frac{\bm{\Pi}_{a}(\hat{\mathbf{n}})}{\omega^{2}+|\mathbf{p}|^{2}s_{a}(\hat{\mathbf{n}})}$, (41)
@@ 3 One-loop diagrams calculation
In this section we calculate all one-loop diagrams necessary to extraction of beta functions. First let us introduce two “master integrals” which are needed for mentioned purposes. So far having the propagator in the form (34) and considering the interactions (12), we need integrals for tadpole and bubble diagrams involving different field species. First let us consider tadpole master integral and then the bubble integral and workout the final answers for pole necessary to calculate beta-function expressed via data like $\mathbf{Z_{t}},\mathbf{Y}$, $\mathbf{C_{\phi,\chi}}$. [^5] In pure dimensional regularization, massless vacuum integrals at zero external momentum are scaleless and vanish, mixing UV and IR. To extract the UV MS pole parts unambiguously we introduce an auxiliary IR regulator $\delta_{\rm IR}>0$; the $1/\varepsilon$ poles are UV-local and independent of $\delta_{\rm IR}$.
@@
In dimensional regularization with $D=4-\varepsilon$ one has $d=3-\varepsilon$. The $1/\varepsilon$ poles in the master integrals come from the radial $|\mathbf{p}|$ integration (e.g. $\int_{0}^{\infty}dx\,x^{-1-\varepsilon/2}\sim 1/\varepsilon$), while the angular averages over $S^{d-1}$ remain finite as $\varepsilon\to 0$ under our positivity assumptions. Therefore, for MS pole parts we may evaluate all angular averages at $d=3$ (i.e. over $S^{2}$); keeping the exact $S^{d-1}$ measure affects only finite terms. Hence, we arrive to
@@
$[\mathbf{T}]_{\text{pole}}=-\frac{1}{4\pi^{2}\epsilon}\left[\left\langle\sum_{a,b}\mathcal{K}(s_{a},s_{b})\mathbf{Z}_{t}^{-1/2}\mathbf{\Pi}_{a}\mathbf{L}\mathbf{\Pi}_{b}\mathbf{Z}_{t}^{-1/2}\right\rangle_{\hat{\mathbf{n}}}\right$], (48)
@@
$\mathbb{B}_{ij;kl}(k)\equiv\int\frac{d\omega}{2\pi}\frac{d^{d}\mathbf{p}}{(2\pi)^{d}}\mathbf{G}_{ik}(p)\mathbf{G}_{jl}(p+k)$
@@
$\mathbb{B}_{ij;kl}(0)\equiv\int\frac{d\omega}{2\pi}\frac{d^{d}\mathbf{p}}{(2\pi)^{d}}\,\mathbf{G}_{ik}(p)\,\mathbf{G}_{jl}(p)$
@@
$\mathbb{B}_{ij;kl}(0)=\frac{\Omega_{d-1}}{2(2\pi)^{d}}\int_{0}^{\infty}dx\;x^{\frac{d}{2}-1}\Big\langle\sum_{a,b}\frac{\tilde{\mathbf{P}}_{a}(\hat{\mathbf{n}},x)_{ik}\tilde{\mathbf{P}}_{b}(\hat{\mathbf{n}},x)_{jl}}{2\sqrt{\kappa_{a}(\hat{\mathbf{n}},x)\kappa_{b}(\hat{\mathbf{n}},x)}\Big(\sqrt{\kappa_{a}(\hat{\mathbf{n}},x)}+\sqrt{\kappa_{b}(\hat{\mathbf{n}},x)}\Big)}\Big\rangle_{\hat{\mathbf{n}}}$.
Performing the expansion at large momentum $|\mathbf{p}|$, one arrives to the following answer for the divergent part of bubble
$[\mathbb{B}_{ij;kl}]_{\text{pole}}=\frac{1}{4\pi^{2}\epsilon}\Big\langle\sum_{a,b}\mathcal{K}(s_{a},s_{b})\mathbf{P}_{a}(\hat{\mathbf{n}})_{ik}\mathbf{P}_{b}(\hat{\mathbf{n}})_{jl}\Big\rangle_{\hat{\mathbf{n}}}$, (52)
@@
Finally let us express all divergent parts of diagrams in compact notation via tadpole and bubble master integral contributions. Diagrams relevant for our calculation are depicted at are well known in Fig.1.
Figure 1: All possible one loop divergent diagrams for the model with the interaction lagrangian (12). The notations $\lambda$ and $h$ are just schematic and we specify couplings for each concrete diagram in the text. Since there are $\delta j_{i}\phi_{i}$ terms in the lagrangian (12), we add the very bottom diagram with one external leg, one propagator and $h_{ijk}$-type vertex as well.
@@ 4 General expressions for beta functions
At fixed bare couplings beta function are defined
@@
where $d_{g}$ is the canonical mass dimension (in $D=4-\varepsilon$) of the dimensionless coupling $g$ in our normalization, fixed by the explicit powers of $\mu$ in the interaction Lagrangian (13). For the couplings in (12) this means
@@
in agreement with the counterterm Lagrangian (14). (Equivalently, one may trade $M^{2}$ for the dimensionless mass matrix $m^{2}\equiv M^{2}/\mu^{2}$ later on.)
@@
$C_{i}$ $=\Big[\mu^{-\frac{D+2}{2}}\,\Gamma^{(1)}_{i,\;{\rm 1\mbox{-}loop}}\Big]_{\rm pole}=-\,\Big[\mu^{2-D}\Delta\Gamma^{(1)}_{i}\Big]_{\rm pole}.$
then the expressions for beta functions has the form
@@
$\beta_{\lambda_{ijkl}}=-\varepsilon\,\lambda_{ijkl}+\frac{1}{2}\Big(\lambda_{ijmn}\,\mathcal{B}_{mn;pq}\,\lambda_{klpq}+\lambda_{ikmn}\,\mathcal{B}_{mn;pq}\,\lambda_{jlpq}+\lambda_{ilmn}\,\mathcal{B}_{mn;pq}\,\lambda_{jkpq}\Big).$ (73)
@@
The cubic couplings renormalize only in the presence of quartic interactions. At one loop the flow is linear in $h$, reflecting the absence of purely cubic one–loop diagrams. From the three–point function pole computed in Sec. 3 one finds
$\beta_{h_{ijk}}=-\Bigl(1+\frac{\varepsilon}{2}\Bigr)h_{ijk}+\frac{1}{2}\Big(\lambda_{ijmn}\,\mathcal{B}_{mn;pq}\,h_{kpq}+\lambda_{ikmn}\,\mathcal{B}_{mn;pq}\,h_{jpq}+\lambda_{jkmn}\,\mathcal{B}_{mn;pq}\,h_{ipq}\Big).$ (74)
@@
where $\widetilde{\mathcal{T}}$ denotes the dimensionless tadpole residue. In particular, for $h\neq 0$ the surface $m^{2}=0$ is not preserved by the RG flow.
@@ 5 Explicit form of beta functions, fixed points and physical interpretation
5.1 Cross-gradients turned off $\mathbf{Y}=0$
@@
Within this limit the field-space projectors are constant, so all direction-dependence collapses into three scalar weights:
$J_{11}=\left\langle\left[\hat{n}^{\top}\mathbf{C}_{1}\hat{n}\right]^{-3/2}\right\rangle_{S^{2}},\quad J_{22}=\left\langle\left[\hat{n}^{\top}\mathbf{C}_{2}\hat{n}\right]^{-3/2}\right\rangle_{S^{2}}$ (80)
$J_{12}=\int_{0}^{1}\left(\operatorname{det}\left[(1-\tau)\mathbf{C}_{1}+\tau\mathbf{C}_{2}\right]\right)^{-1/2}d\tau$
@@
$=\frac{1}{8\pi^{2}}\Big(J_{11}\delta_{m1}\delta_{n1}\delta_{p1}\delta_{q1}+J_{22}\delta_{m2}\delta_{n2}\delta_{p2}\delta_{q2}+J_{12}(\delta_{m1}\delta_{n2}\delta_{p1}\delta_{q2}+\delta_{m2}\delta_{n1}\delta_{p2}\delta_{q1})\Big).$ (81)
@@
$\mathcal{T}\equiv\varepsilon\,\mathbf{T}=-\frac{1}{8\pi^{2}}\begin{pmatrix}J_{11}M_{11}^{2}&J_{12}M_{12}^{2}\\ J_{12}M_{12}^{2}&J_{22}M_{22}^{2}\end{pmatrix}\qquad(\mathbf{Z}_{t}=\mathbf{1},\;\mathbf{Y}=0).$ (82)
@@
After some algebra one can obtain the explicit form of $J_{ij}$ in terms of eigenvalues data of $\mathbf{C_{a}}$. In dimensional regularization with $D=4-\epsilon$ one has $d=3-\epsilon$. The MS pole coefficients require the angular averages only at $\epsilon=0$, so throughout this subsection we evaluate the weights at $d=3$, i.e. over $S^{2}$. (Equivalently, one may keep $S^{d-1}$ and set $\epsilon\to 0$ at the end; the difference contributes only to finite parts.)
@@
$J_{12}=\left(\operatorname{det}\mathbf{C}_{1}\right)^{-1/2}\int_{0}^{1}\prod_{i=1}^{3}\left((1-\tau)+\tau\gamma_{i}\right)^{-1/2}d\tau$
@@
$J_{12}^{2}\leq J_{11}J_{22}.$ (84)
In these notations we obtain that beta-functions can be written down explicitly in the form resembling the well-known answer for two-fields beta-function, but weighted by $J_{ij}$ in appropriate places. Namely for $\lambda_{i}$ set of constants we obtain
$\beta_{\lambda_{1}}=-\varepsilon\lambda_{1}+\frac{3}{16\pi^{2}}\left[J_{11}\lambda_{1}^{2}+2J_{12}\lambda_{4}^{2}+J_{22}\lambda_{3}^{2}\right],$ (85a)
$\beta_{\lambda_{2}}=-\varepsilon\lambda_{2}+\frac{3}{16\pi^{2}}\left[J_{22}\lambda_{2}^{2}+2J_{12}\lambda_{5}^{2}+J_{11}\lambda_{3}^{2}\right],$ (85b)
$\beta_{\lambda_{3}}=-\varepsilon\lambda_{3}$
@@
$\beta_{\lambda_{4}}=-\varepsilon\lambda_{4}+\frac{3}{16\pi^{2}}\left[J_{11}\lambda_{1}\lambda_{4}+2J_{12}\lambda_{3}\lambda_{4}+J_{22}\lambda_{3}\lambda_{5}\right],$ (85d)
$\beta_{\lambda_{5}}=-\varepsilon\lambda_{5}+\frac{3}{16\pi^{2}}\left[J_{22}\lambda_{2}\lambda_{5}+2J_{12}\lambda_{3}\lambda_{5}+J_{11}\lambda_{3}\lambda_{4}\right],$ (85e)
@@
$\beta_{h_{1}}=-\left(1+\frac{\varepsilon}{2}\right)h_{1}+\frac{3}{16\pi^{2}}\left[J_{11}\lambda_{1}h_{1}+2J_{12}\lambda_{4}h_{3}+J_{22}\lambda_{3}h_{4}\right],$ (86a)
$\beta_{h_{2}}=-\left(1+\frac{\varepsilon}{2}\right)h_{2}+\frac{3}{16\pi^{2}}\left[J_{22}\lambda_{2}h_{2}+2J_{12}\lambda_{5}h_{4}+J_{11}\lambda_{3}h_{3}\right],$ (86b)
$\beta_{h_{3}}=-\left(1+\frac{\varepsilon}{2}\right)h_{3}$
@@
$\beta_{h_{4}}=-\left(1+\frac{\varepsilon}{2}\right)h_{4}$
@@
$\beta_{j_{\phi}}$ $=-\left(3-\frac{\varepsilon}{2}\right)j_{\phi}+\frac{1}{16\pi^{2}}\Big(J_{11}\,h_{1}\,m_{11}^{2}+2J_{12}\,h_{3}\,m_{12}^{2}+J_{22}\,h_{4}\,m_{22}^{2}\Big),$ (88a)
$\beta_{j_{\chi}}$ $=-\left(3-\frac{\varepsilon}{2}\right)j_{\chi}+\frac{1}{16\pi^{2}}\Big(J_{11}\,h_{3}\,m_{11}^{2}+2J_{12}\,h_{4}\,m_{12}^{2}+J_{22}\,h_{2}\,m_{22}^{2}\Big),$ (88b)
@@
$\beta_{\lambda_{1}}$ $=-\varepsilon\lambda_{1}+\frac{3}{16\pi^{2}}\left[J_{11}\lambda_{1}^{2}+J_{22}\lambda_{3}^{2}\right],$
$\beta_{\lambda_{2}}$ $=-\varepsilon\lambda_{2}+\frac{3}{16\pi^{2}}\left[J_{22}\lambda_{2}^{2}+J_{11}\lambda_{3}^{2}\right],$
$\beta_{\lambda_{3}}$ $=-\varepsilon\lambda_{3}+\frac{1}{16\pi^{2}}\left[J_{11}\lambda_{1}\lambda_{3}+4J_{12}\lambda_{3}^{2}+J_{22}\lambda_{2}\lambda_{3}\right],$
and formal solutions to set equations $\beta_{\lambda_{i}}=0$ corresponds to different fixed points of our model. Namely, we have two solution we call $S_{\pm}$ (corresponding to $\pm$ sign choices):
$\lambda_{1}^{*}=\frac{\frac{16\pi^{2}}{3}\varepsilon\cdot 3-4J_{12}\lambda_{3}^{*}}{2J_{11}},$ (89a)
$\lambda_{2}^{*}=\frac{\frac{16\pi^{2}}{3}\varepsilon\cdot 3-4J_{12}\lambda_{3}^{*}}{2J_{22}},$ (89b)
$\lambda_{3}^{*}=\frac{\frac{16\pi^{2}}{3}\varepsilon\left(4J_{12}\pm\sqrt{4J_{12}^{2}-3J_{11}J_{22}}\right)}{2\left(4J_{12}^{2}+J_{11}J_{22}\right)},$ (89c)
which is real-valued only when $4J_{12}^{2}\geq 3J_{11}J_{22}$. Besides the coupled solutions $S_{\pm}$ with $\lambda_{3}^{\star}\neq 0$, the equations $\beta_{\lambda_{1}}=\beta_{\lambda_{2}}=\beta_{\lambda_{3}}=0$ also admit the standard fixed points:
  • Gaussian: $(\lambda_{1}^{\star},\lambda_{2}^{\star},\lambda_{3}^{\star})=(0,0,0)$.
  • Single-field Wilson–Fisher points (with $\lambda_{3}^{\star}=0$):
    $(\lambda_{1}^{\star},\lambda_{2}^{\star},\lambda_{3}^{\star})=\Big(\frac{16\pi^{2}}{3}\frac{\varepsilon}{J_{11}},\,0,\,0\Big),\qquad(\lambda_{1}^{\star},\lambda_{2}^{\star},\lambda_{3}^{\star})=\Big(0,\,\frac{16\pi^{2}}{3}\frac{\varepsilon}{J_{22}},\,0\Big).$
  • Decoupled two-field Wilson–Fisher point (with $\lambda_{3}^{\star}=0$):
    $(\lambda_{1}^{\star},\lambda_{2}^{\star},\lambda_{3}^{\star})=\Big(\frac{16\pi^{2}}{3}\frac{\varepsilon}{J_{11}},\,\frac{16\pi^{2}}{3}\frac{\varepsilon}{J_{22}},\,0\Big).$
We focus below on the coupled fixed points with $\lambda_{3}^{\star}\neq 0$.
@@
while in general it belongs to $\rho\in(0,1]$ Solution corresponding to $S_{-}$ is the IR-stable attractor generalizing ordinary Wilson-Fisher fixed point and reducing to it when all velocities are set to unity and for appropriate choice of constants. One can observe the effect of destruction of WF fixed point corresponding to (89) due to existence of restrictions $4J_{12}^{2}>3J_{11}J_{22}$. The solution $S_{+}$ is the IR fixed point of saddle type containing one repulsive direction.

### v2

@@ 2 Setup: theory and its decomposition
The resolution for $(\omega^{2}\mathbf{1}+|\mathbf{p}|^{2}\mathbf{S}(\hat{\mathbf{n}}))^{-1}$ then reads
$(\omega^{2}\mathbf{1}+|\mathbf{p}|^{2}\mathbf{S}(\hat{\mathbf{n}}))^{-1}=\sum_{a=1}^{2}\frac{\bm{\Pi}_{a}(\hat{\mathbf{n}})}{\omega^{2}+|\mathbf{p}|^{2}s_{a}(\hat{\mathbf{n}})}$, (39)
@@ 3 One-loop diagrams calculation
In this section we calculate all one-loop diagrams necessary for the extraction of beta functions. First let us introduce two “master integrals” which are needed for mentioned purposes. So far having the propagator in the form (32) and considering the interactions (12), we need integrals for tadpole and bubble diagrams involving different field species. First let us consider tadpole master integral and then the bubble integral and work out the final answers for pole necessary to calculate beta function expressed via data like $\mathbf{Z}_{t},\mathbf{Y}$, $\mathbf{C_{\phi,\chi}}$. [^6] In pure dimensional regularization, massless vacuum integrals at zero external momentum are scaleless and vanish, mixing UV and IR. To extract the UV MS pole parts unambiguously we introduce an auxiliary IR regulator $\delta_{\rm IR}>0$; the $1/\varepsilon$ poles are UV-local and independent of $\delta_{\rm IR}$.
@@
In dimensional regularization with $D=4-\varepsilon$ one has $d=3-\varepsilon$. The $1/\varepsilon$ poles in the master integrals come from the radial $|\mathbf{p}|$ integration (e.g. $\int_{0}^{\infty}dx\,x^{-1-\varepsilon/2}\sim 1/\varepsilon$), while the angular averages over $S^{d-1}$ remain finite as $\varepsilon\to 0$ under our positivity assumptions. Therefore, for MS pole parts we may evaluate all angular averages at $d=3$ (i.e. over $S^{2}$); keeping the exact $S^{d-1}$ measure affects only finite terms. Hence, we arrive at
@@
$\mathbb{T}\equiv[\mathbf{T}]_{\text{pole}}=-\frac{1}{4\pi^{2}\epsilon}\left[\left\langle\sum_{a,b}\mathcal{K}(s_{a},s_{b})\mathbf{Z}_{t}^{-1/2}\mathbf{\Pi}_{a}\mathbf{L}\mathbf{\Pi}_{b}\mathbf{Z}_{t}^{-1/2}\right\rangle_{\hat{\mathbf{n}}}\right$], (46)
@@
$\mathbf{B}_{ij;kl}(k)\equiv\int\frac{d\omega}{2\pi}\frac{d^{d}\mathbf{p}}{(2\pi)^{d}}\mathbf{G}_{ik}(p)\mathbf{G}_{jl}(p+k)$
@@
$\mathbf{B}_{ij;kl}(0)\equiv\int\frac{d\omega}{2\pi}\frac{d^{d}\mathbf{p}}{(2\pi)^{d}}\,\mathbf{G}_{ik}(p)\,\mathbf{G}_{jl}(p)$
@@
$\mathbf{B}_{ij;kl}(0)=\frac{\Omega_{d-1}}{2(2\pi)^{d}}\int_{0}^{\infty}dx\;x^{\frac{d}{2}-1}\Big\langle\sum_{a,b}\frac{\tilde{\mathbf{P}}_{a}(\hat{\mathbf{n}},x)_{ik}\tilde{\mathbf{P}}_{b}(\hat{\mathbf{n}},x)_{jl}}{2\sqrt{\kappa_{a}(\hat{\mathbf{n}},x)\kappa_{b}(\hat{\mathbf{n}},x)}\Big(\sqrt{\kappa_{a}(\hat{\mathbf{n}},x)}+\sqrt{\kappa_{b}(\hat{\mathbf{n}},x)}\Big)}\Big\rangle_{\hat{\mathbf{n}}}$.
Performing the expansion at large momentum $|\mathbf{p}|$ after some algebra one arrives to the following answer for the divergent part of bubble
$\mathbb{B}_{ij;kl}\equiv[\mathbf{B}_{ij;kl}]_{\text{pole}}=\frac{1}{4\pi^{2}\epsilon}\Big\langle\sum_{a,b}\mathcal{K}(s_{a},s_{b})\mathbf{P}_{a}(\hat{\mathbf{n}})_{ik}\mathbf{P}_{b}(\hat{\mathbf{n}})_{jl}\Big\rangle_{\hat{\mathbf{n}}}$, (50)
@@
Finally let us express all divergent parts of diagrams in compact notation via tadpole and bubble master integral contributions. Diagrams relevant for our calculation are well known and depicted in Fig.1.
Figure 1: All possible one loop divergent diagrams for the model with the interaction Lagrangian (12). The notations $\lambda$ and $h$ are just schematic and we specify couplings for each concrete diagram in the text. Since there are $\delta j_{i}\phi_{i}$ terms in the Lagrangian (12), we add the bottom diagram with one external leg, one propagator and $h_{ijk}$-type vertex as well.
@@ 4 General expressions for beta functions
At fixed bare couplings beta functions are defined
@@
where $d_{g}$ is the canonical mass dimension (in $D=4-\varepsilon$) of the bare coupling $g_{0}$ in our normalization, fixed by the explicit powers of $\mu$ in the interaction Lagrangian (13). For the couplings in (12) this means
@@
in agreement with the counterterm Lagrangian (2).
@@
$C_{i}$ $=\Big[\mu^{-\frac{D+2}{2}}\,\Gamma^{(1)}_{i,\;{\rm 1\mbox{-}loop}}\Big]_{\rm pole}=-\,\Big[\mu^{2-D}\Delta\Gamma^{(1)}_{i}\Big]_{\rm pole},$
then the expressions for beta functions have the form
@@
$\beta_{\lambda_{ijkl}}=-\varepsilon\,\lambda_{ijkl}$ (71)
$+\frac{1}{2}\Big(\lambda_{ijmn}\,\mathcal{B}_{mn;pq}\,\lambda_{klpq}+\lambda_{ikmn}\,\mathcal{B}_{mn;pq}\,\lambda_{jlpq}+\lambda_{ilmn}\,\mathcal{B}_{mn;pq}\,\lambda_{jkpq}\Big).$
@@
The cubic couplings renormalize only in the presence of quartic interactions. At one loop the flow is linear in $h$, reflecting the absence of UV-divergent purely cubic one–loop diagrams near $D=4$. From the three–point function pole computed in Sec. 3 one finds
$\beta_{h_{ijk}}=-\Bigl(1+\frac{\varepsilon}{2}\Bigr)h_{ijk}$ (72)
$+\frac{1}{2}\Big(\lambda_{ijmn}\,\mathcal{B}_{mn;pq}\,h_{kpq}+\lambda_{ikmn}\,\mathcal{B}_{mn;pq}\,h_{jpq}+\lambda_{jkmn}\,\mathcal{B}_{mn;pq}\,h_{ipq}\Big).$
@@
where $\widetilde{\mathcal{T}}_{ij}\equiv\frac{\mathcal{T}_{ij}}{\mu^{2}}$ denotes the dimensionless tadpole residue, while $B_{mn;pq}$ is already dimensionless. In particular, for $h\neq 0$ the surface $m^{2}=0$ is not preserved by the RG flow.
@@ 5 Explicit form of beta functions, fixed points and physical interpretation
5.1 Cross-gradients turned off: $\mathbf{Y}=0$
@@
$J_{11}=\left\langle\left[\hat{n}^{\top}\mathbf{C}_{1}\hat{n}\right]^{-3/2}\right\rangle_{S^{2}},\quad J_{22}=\left\langle\left[\hat{n}^{\top}\mathbf{C}_{2}\hat{n}\right]^{-3/2}\right\rangle_{S^{2}},$ (78)
$J_{12}=\int_{0}^{1}\left(\operatorname{det}\left[(1-\tau)\mathbf{C}_{1}+\tau\mathbf{C}_{2}\right]\right)^{-1/2}d\tau,$
@@
$=\frac{1}{8\pi^{2}}\Big(J_{11}\delta_{m1}\delta_{n1}\delta_{p1}\delta_{q1}+J_{22}\delta_{m2}\delta_{n2}\delta_{p2}\delta_{q2}+J_{12}(\delta_{m1}\delta_{n2}\delta_{p1}\delta_{q2}+\delta_{m2}\delta_{n1}\delta_{p2}\delta_{q1})\Big),$ (79)
@@
$\mathcal{T}\equiv\varepsilon\,\mathbf{T}=-\frac{1}{8\pi^{2}}\begin{pmatrix}J_{11}M_{11}^{2}&J_{12}M_{12}^{2}\\ J_{12}M_{12}^{2}&J_{22}M_{22}^{2}\end{pmatrix}$. (80)
@@
After some algebra one can obtain the explicit form of $J_{ij}$ in terms of eigenvalues data of $\mathbf{C_{a}}$. In dimensional regularization with $d=3-\epsilon$, the MS pole coefficients require the angular averages only at $\epsilon=0$ (equivalently, one may keep $S^{d-1}$ and set $\epsilon\to 0$ at the end; the difference contributes only to finite parts).
@@
$J_{12}=\left(\operatorname{det}\mathbf{C}_{1}\right)^{-1/2}\int_{0}^{1}\prod_{i=1}^{3}\left((1-\tau)+\tau\gamma_{i}\right)^{-1/2}d\tau,$
@@
$J_{12}^{2}\leq J_{11}J_{22},$ (82)
and that is why $\rho\in(0,1]$; the derivation of this (82) inequality is presented in App. B.2. In these notations we obtain that beta functions can be written down explicitly in the form resembling the well-known answer for two-fields beta function, but weighted by $J_{ij}$ in appropriate places. Namely, for $\lambda_{i}$ set of constants we obtain
$\beta_{\lambda_{1}}=-\epsilon\lambda_{1}+\frac{3}{16\pi^{2}}\left[J_{11}\lambda_{1}^{2}+2J_{12}\lambda_{4}^{2}+J_{22}\lambda_{3}^{2}\right],$ (83a)
$\beta_{\lambda_{2}}=-\epsilon\lambda_{2}+\frac{3}{16\pi^{2}}\left[J_{22}\lambda_{2}^{2}+2J_{12}\lambda_{5}^{2}+J_{11}\lambda_{3}^{2}\right],$ (83b)
$\beta_{\lambda_{3}}=-\epsilon\lambda_{3}$
@@
$\beta_{\lambda_{4}}=-\epsilon\lambda_{4}+\frac{3}{16\pi^{2}}\left[J_{11}\lambda_{1}\lambda_{4}+2J_{12}\lambda_{3}\lambda_{4}+J_{22}\lambda_{3}\lambda_{5}\right],$ (83d)
$\beta_{\lambda_{5}}=-\epsilon\lambda_{5}+\frac{3}{16\pi^{2}}\left[J_{22}\lambda_{2}\lambda_{5}+2J_{12}\lambda_{3}\lambda_{5}+J_{11}\lambda_{3}\lambda_{4}\right],$ (83e)
@@
$\beta_{h_{1}}=-\left(1+\frac{\epsilon}{2}\right)h_{1}+\frac{3}{16\pi^{2}}\left[J_{11}\lambda_{1}h_{1}+2J_{12}\lambda_{4}h_{3}+J_{22}\lambda_{3}h_{4}\right],$ (84a)
$\beta_{h_{2}}=-\left(1+\frac{\epsilon}{2}\right)h_{2}+\frac{3}{16\pi^{2}}\left[J_{22}\lambda_{2}h_{2}+2J_{12}\lambda_{5}h_{4}+J_{11}\lambda_{3}h_{3}\right],$ (84b)
$\beta_{h_{3}}=-\left(1+\frac{\epsilon}{2}\right)h_{3}$
@@
$\beta_{h_{4}}=-\left(1+\frac{\epsilon}{2}\right)h_{4}$
@@
$\beta_{j_{\phi}}$ $=-\left(3-\frac{\epsilon}{2}\right)j_{\phi}+\frac{1}{16\pi^{2}}\Big(J_{11}\,h_{1}\,m_{11}^{2}+2J_{12}\,h_{3}\,m_{12}^{2}+J_{22}\,h_{4}\,m_{22}^{2}\Big),$ (86a)
$\beta_{j_{\chi}}$ $=-\left(3-\frac{\epsilon}{2}\right)j_{\chi}+\frac{1}{16\pi^{2}}\Big(J_{11}\,h_{3}\,m_{11}^{2}+2J_{12}\,h_{4}\,m_{12}^{2}+J_{22}\,h_{2}\,m_{22}^{2}\Big),$ (86b)
@@
$\beta_{\lambda_{1}}$ $=-\epsilon\lambda_{1}+\frac{3}{16\pi^{2}}\left[J_{11}\lambda_{1}^{2}+J_{22}\lambda_{3}^{2}\right],$
$\beta_{\lambda_{2}}$ $=-\epsilon\lambda_{2}+\frac{3}{16\pi^{2}}\left[J_{22}\lambda_{2}^{2}+J_{11}\lambda_{3}^{2}\right],$
$\beta_{\lambda_{3}}$ $=-\epsilon\lambda_{3}+\frac{1}{16\pi^{2}}\left[J_{11}\lambda_{1}\lambda_{3}+4J_{12}\lambda_{3}^{2}+J_{22}\lambda_{2}\lambda_{3}\right],$
and formal solutions to set equations $\beta_{\lambda_{i}}=0$ corresponds to different fixed points of our model. Namely, we have two solutions we call $S_{\pm}$ (corresponding to $\pm$ sign choices):
$\lambda_{1}^{*}=\frac{\frac{16\pi^{2}}{3}\epsilon\cdot 3-4J_{12}\lambda_{3}^{*}}{2J_{11}},$ (87a)
$\lambda_{2}^{*}=\frac{\frac{16\pi^{2}}{3}\epsilon\cdot 3-4J_{12}\lambda_{3}^{*}}{2J_{22}},$ (87b)
$\lambda_{3}^{*}=\frac{\frac{16\pi^{2}}{3}\epsilon\left(4J_{12}\pm\sqrt{4J_{12}^{2}-3J_{11}J_{22}}\right)}{2\left(4J_{12}^{2}+J_{11}J_{22}\right)},$ (87c)
which are real-valued only when $4J_{12}^{2}\geq 3J_{11}J_{22}$. Besides the coupled solutions $S_{\pm}$ with $\lambda_{3}^{*}\neq 0$, the equations $\beta_{\lambda_{1}}=\beta_{\lambda_{2}}=\beta_{\lambda_{3}}=0$ also admit the standard fixed points:
  • Gaussian: $(\lambda_{1}^{*},\lambda_{2}^{*},\lambda_{3}^{*})=(0,0,0)$.
  • Single-field Wilson–Fisher points (with $\lambda_{3}^{*}=0$):
    $(\lambda_{1}^{*},\lambda_{2}^{*},\lambda_{3}^{*})=\Big(\frac{16\pi^{2}}{3}\frac{\varepsilon}{J_{11}},\,0,\,0\Big),\qquad(\lambda_{1}^{*},\lambda_{2}^{*},\lambda_{3}^{*})=\Big(0,\,\frac{16\pi^{2}}{3}\frac{\varepsilon}{J_{22}},\,0\Big).$
  • Decoupled two-field Wilson–Fisher point (with $\lambda_{3}^{*}=0$):
    $(\lambda_{1}^{*},\lambda_{2}^{*},\lambda_{3}^{*})=\Big(\frac{16\pi^{2}}{3}\frac{\varepsilon}{J_{11}},\,\frac{16\pi^{2}}{3}\frac{\varepsilon}{J_{22}},\,0\Big).$
We focus below on the coupled fixed points with $\lambda_{3}^{*}\neq 0$.
@@
while in general it belongs to $\rho\in(0,1]$. Solution corresponding to $S_{-}$ is the IR-stable attractor generalizing ordinary Wilson-Fisher fixed point and reducing to it when all velocities are set to unity and for appropriate choice of constants. The coupled fixed point exists only when the mismatch is sufficiently small, $\rho\geq\sqrt{3}/2$. When $\rho<\sqrt{3}/2$ the coupled solution becomes complex and disappears from the real coupling space. The solution $S_{+}$ is the IR fixed point of saddle type containing one repulsive direction.